### v2

# Localization of spectral Turán-type theorems

Abstract
Let $G$ be a graph, and let $v$ and $e$ be a vertex and an edge of $G$, respectively. Define $c(v)$ (resp. $c(e)$) to be the order of the largest clique in $G$ containing $v$ (resp. $e$). Denote the adjacency eigenvalues of $G$ by $\lambda_{1}\geq\cdots\geq\lambda_{n}$. We study localized refinements of spectral Turán-type theorems by replacing global parameters such as the clique number $\omega(G)$, size $m$ and order $n$ of $G$ with local quantities $c(v)$ and $c(e)$.
Motivated by a conjecture of Elphick, Linz and Wocjan (2024), we first propose a vertex-localized strengthening of Wilf’s inequality:
$\sqrt{s^{+}(G)}\leq\sum_{v\in V(G)}\left(1-\frac{1}{c(v)}\right),$
where $s^{+}(G)=\sum_{\lambda_{i}>0}\lambda_{i}^{2}$. Inspired by the Bollobás-Nikiforov conjecture (2007) on the first two eigenvalues, we then introduce an edge-localized analogue:
$\lambda_{1}^{2}(G)+\lambda_{2}^{2}(G)\leq\sum_{e\in E(G)}2\left(1-\frac{1}{c(e)}\right).$
As evidence of their validity, we verify the above conjectures for diamond-free graphs and random graphs. We also propose strengthening of the spectral versions of the Erdős-Stone-Simonovits Theorem by replacing the spectral radius with $\sqrt{s^{+}(G)}$ and establish it for all $F$-free graphs with $\chi(F)=3$. A key ingredient in our proofs is a general upper bound relating $\sqrt{s^{+}(G)}$ to the triangle count $t(G)$. Finally, we prove a localized version of Nikiforov’s walk inequality and conjecture a stronger localized version. These results contribute to the broader program of localizing spectral extremal inequalities.
Keywords: Localization, Tuŕan-type theorems, Spectral radius, Second largest eigenvalue, Square energy, Clique number.
MSC: 05C50, 05C35, 15A42.
1 Introduction
1.1 Notation
We use standard graph theory notation and terminology throughout the paper. We assume all graphs are simple and connected unless specified otherwise. Let $G=(V(G),E(G))$ be a graph of order $n$ and size $m$. Let $t(G)$, $\Delta(G)$, $\omega(G)$ and $\chi(G)$ denote the number of triangles, maximum degree, clique number and chromatic number of $G$, respectively. For a graph $F$, we say that $G$ is $F$-free if $G$ does not contain $F$ as a subgraph (not necessarily induced). Let $\deg(v)$ denote the degree of vertex $v$ in $G$. For an edge $e\in E(G)$ (or a vertex $v\in V(G)$), we will denote by $c(e)$ (resp. $c(v)$) the order of the largest clique in $G$ containing the edge $e$ (resp. the vertex $v$). The adjacency matrix of $G$ is an $n\times n$ matrix $A(G)=[a_{uv}]$, where $a_{uv}=1$ if $u$ and $v$ are adjacent, and 0 otherwise. The eigenvalues of $G$ are the eigenvalues of $A(G)$. Since $A(G)$ is a real symmetric matrix, all eigenvalues of $A(G)$ are real and can be listed as $\lambda_{1}(G)\geq\cdots\geq\lambda_{n}(G)$. Let
[…]
We call $s^{+}(G)$ (resp. $s^{-}(G)$) the positive (resp. negative) square energy of $G$. Let $n^{+}$ and $n^{-}$ denote the number of positive and negative eigenvalues of $A(G)$, respectively.
For a graph $G$, we can assign a (non-negative) weight function $w:E(G)\rightarrow\mathbb{R}^{\geq 0}$ to its edges and define a weighted adjacency matrix $W(G)=[w_{uv}]$, where $w_{uv}=w(uv)$ if $u$ and $v$ are adjacent, and $w_{uv}=0$ otherwise. For an integer $r\geq 1$ and a vertex $v\in V(G)$, denote by $w_{r}(v)$ the number of walks with $r$ vertices starting from $v$ in $G$. The total number of $r$-walks in $G$ is then given by
$w_{r}(G)=\sum_{v\in V(G)}w_{r}(v).$
1.2 Known results
One of the central themes in graph theory is understanding how local constraints influence global structure. A classic example is determining the maximum number of edges a graph can have while avoiding a specific forbidden subgraph. This area of study, known as extremal graph theory, seeks to characterize such optimal graphs under given restrictions. A foundational result in this field is Tuŕan’s Theorem, which provides an exact bound on the number of edges in a graph that does not contain a complete subgraph $K_{\omega+1}$ .
Theorem 1.1 (Tuŕan’s Theorem).
For any graph $G$ with the clique number $\omega$, we have
$m\leq\frac{n^{2}}{2}\left(1-\frac{1}{\omega}\right).$
Equality holds if and only if $G$ is a complete regular $\omega$-partite graph.
Wilf [35] obtained the following spectral generalization of Tuŕan’s result.
Theorem 1.2 (Wilf’s Inequality [35]).
For any graph $G$, we have
$\lambda_{1}(G)\leq n\left(1-\frac{1}{\omega}\right).$
Equality holds if and only if $G$ is a complete regular $\omega$-partite graph.
Wilf’s Inequality is stronger than Tuŕan’s Theorem since the spectral radius is always larger than the average degree of the graph, i.e., $\lambda_{1}\geq\frac{2m}{n}$.
Furthermore, Edwards and Elphick [9] conjectured the following generalization, which was later proved by Nikiforov [30], now known as the spectral Tuŕan’s Theorem.
Theorem 1.3 (Spectral Tuŕan’s Theorem [30]).
For any graph $G$, we have
$\lambda_{1}^{2}(G)\leq 2m\left(1-\frac{1}{\omega}\right).$
The following extension of Turán’s Theorem to $F$-free graphs is well-known.
Theorem 1.4 (Erdős-Stone-Simonovits Theorem [13, 14]).
For any $F$-free graph $G$ where $\chi(F)\geq 3$, we have
$m\leq\frac{n^{2}}{2}\left(1-\frac{1}{\chi(F)-1}+o(1)\right).$
Nikiforov [31] established the spectral version of the above theorem in the spirit of Wilf’s Inequality.
Theorem 1.5 ([31]).
For any $F$-free graph $G$ where $\chi(F)\geq 3$, we have
$\lambda_{1}(G)\leq n\left(1-\frac{1}{\chi(F)-1}+o(1)\right).$
Li, Liu and Zhang [21] proved an edge-spectral generalization of Erdős-Stone-Simonovits Theorem.
Theorem 1.6 ([21]).
For any $F$-free graph $G$ where $\chi(F)\geq 3$, we have
$\lambda_{1}^{2}(G)\leq 2m\left(1-\frac{1}{\chi(F)-1}+o(1)\right).$
Recently, there has been an interest in generalizing classical results by assigning a weight function to the edges (resp. vertices) of the graph and then proving a bound on this weight function which implies the classical result. This is called edge localization (resp. vertex localization). Malec and Tompkins [28] edge-localized a wide range of classical results in graph theory, and many interesting follow-up papers have appeared, see [1, 4, 19, 40, 39]. Tuŕan’s Theorem was edge-localized independently by Bradač [7] and Malec and Tompkins [28]. Recently, Adak and Chandran [2] vertex-localized Tuŕan’s Theorem. Liu and Ning [27] obtained the following edge localization of the spectral Tuŕan’s Theorem.
Theorem 1.7 (Edge-localized Spectral Tuŕan’s Theorem [27]).
For any graph $G$, we have
$\lambda_{1}^{2}(G)\leq\sum_{e\in E(G)}2\left(1-\frac{1}{c(e)}\right).$
We note here that in Theorems 1.3 and 1.7, equality holds if and only if $G$ is a complete bipartite graph when $\omega=2$ or a regular complete $\omega$-partite graph when $\omega\geq 3$.
Recently, Liu and Ning generalized Theorem 1.7 to weighted graphs, which implies the results of Bradač, and Malec and Tompkins for Tuŕan’s Theorem.
Theorem 1.8 (Weighted Edge-localized Spectral Tuŕan’s Theorem [26]).
For any connected graph $G$ with weighted adjacency matrix $W(G)$, we have
$\lambda_{1}^{2}(W(G))\leq\sum_{e\in E(G)}2\left(1-\frac{1}{c(e)}\right)w(e)^{2}.$
1.3 Our contribution
Elphick, Linz and Wocjan [11] proposed the following generalization of Wilf’s Inequality.
Conjecture 1.9 ([11]).
For any graph $G$, we have
$\sqrt{s^{+}(G)}\leq n\left(1-\frac{1}{\omega}\right).$
In [11], the authors have verified the above conjecture for weakly-perfect graphs, Kneser graphs, triangle-free graphs and random graphs. We prove a weaker but non-trivial upper bound for $\sqrt{s^{+}}$ of any graph $G$ as follows.
Theorem 1.10.
Let $G$ be a graph of order $n$ and the clique number $\omega$. Then
$\sqrt{s^{+}(G)}\leq n\sqrt{1-\frac{1}{\omega}-\frac{1}{\omega^{2}}}\ .$
Next, we propose a vertex-localized version of the above conjecture.
Conjecture 1.11.
For any graph $G$, we have
$\sqrt{s^{+}(G)}\leq\sum_{v\in V}\left(1-\frac{1}{c(v)}\right).$
We present a case for this conjecture by proving it for diamond-free graphs (Theorem 3.3), and almost all graphs using Erdős-Rényi random graph $G(n,1/2)$ (Theorem 3.5). We have also computationally verified this conjecture for graphs up to 9 vertices here. To prove Conjecture 1.11 for diamond-free graphs, we first establish the following general upper bound for $s^{+}$ in terms of the number of triangles $t(G)$, which we believe is of independent interest.
Theorem 1.12.
Let $G$ be a graph of order $n$, size $m$, and $t(G)$ be the number of triangles in $G$. Then
$\sqrt{s^{+}(G)}\leq\frac{n}{2}+\frac{3t(G)}{\lambda^{2}_{1}(G)}.$
Equality holds if and only if $G$ is a complete bipartite graph.
Contrast this with a known result of Elphick and Linz [10] which states that $\sqrt{s^{-}(G)}\leq\frac{n}{2}$ for any graph $G$ of order $n$. This further suggests an asymmetry between $s^{+}$ and $s^{-}$; also refer [37].
For the $F$-free graphs, we propose a strengthening of Nikiforov’s Theorem 1.5 by replacing $\lambda_{1}$ with $\sqrt{s^{+}}$.
Conjecture 1.13.
For any $F$-free graph $G$ where $\chi(F)\geq 3$, we have
$\sqrt{s^{+}(G)}\leq n\left(1-\frac{1}{\chi(F)-1}+o(1)\right).$
We prove this conjecture for $F$-free graphs when $\chi(F)=3$ (Theorem 3.4).
Bollobás and Nikiforov [6] proposed to improve Theorem 1.3 by introducing $\lambda_{2}$.
Conjecture 1.14 (Bollobás - Nikiforov Conjecture [6]).
For any graph $G\neq K_{n}$, we have
$\lambda_{1}^{2}(G)+\lambda^{2}_{2}(G)\leq 2m\left(1-\frac{1}{\omega}\right).$
The above conjecture has been verified for weakly perfect graphs by Ando and Lin [3], triangle-free graphs by Lin, Ning and Wu [24], regular graphs by Zhang [38], and random graphs by Liu and Bu [25]. Kumar and Pragada [20] verified the conjecture for graphs with few triangles, which includes cycle-free graphs, book-free graphs and planar graphs. Zeng and Zhang [36] proved the conjecture for line graphs and also strengthened some results from [20]. Motivated by the recent localization results, it is only natural to ask if the following edge-localized refinement of the Bollobás-Nikiforov Conjecture is true.
Conjecture 1.15.
For any graph $G\neq K_{n}$, we have
$\lambda_{1}^{2}(G)+\lambda_{2}^{2}(G)\leq\sum_{e\in E(G)}2\left(1-\frac{1}{c(e)}\right).$
Clearly, Conjecture 1.15 is true for triangle-free graphs. As evidence for this conjecture, we prove it for the next non-trivial case of diamond-free graphs (Theorem 4.4) whenever $t(G)\notin\{1,2,3,4\}$, and random graphs (Theorem 4.6). We have computationally verified this conjecture for graphs up to 9 vertices here.
We note that equality in Conjecture 1.15 is attained not just by regular complete multipartite graphs (barring $K_{n}$), but also by the disjoint union of two regular complete multipartite graphs of not necessarily the same regularities. This suggests that Conjecture 1.15 captures a rich extremal phenomenon.
Recently, Li, Liu and Zhang [22] proposed a strengthening of the Bollobás-Nikiforov Conjecture for $F$-free graphs where $F$ is a color-critical graph.
Conjecture 1.16 ([22]).
Let $F$ be a color-critical graph with $\chi(F)\geq 4$. Then any $F$-free graph $G$ with sufficiently large size $m$ satisfies
$\lambda_{1}^{2}(G)+\lambda^{2}_{2}(G)\leq 2m\left(1-\frac{1}{\chi(F)-1}\right).$
We propose a weaker bound, but for any graph $F$ with $\chi(F)\geq 3$.
Conjecture 1.17.
For any $F$-free graph $G$ where $\chi(F)\geq 3$, we have
$\lambda_{1}^{2}(G)+\lambda^{2}_{2}(G)\leq 2m\left(1-\frac{1}{\chi(F)-1}+o(1)\right).$
We prove this statement when $\chi(F)=3$ (Theorem 4.5).
Nikiforov [30] gave an upper bound on $\lambda_{1}$ using the number of $r$-walks in the graph.
Theorem 1.18 (Nikiforov’s Walk Inequality [30]).
Let $G$ be a graph. For any integer $r\geq 1$, we have
$\lambda_{1}^{r}(G)\leq w_{r}(G)\left(1-\frac{1}{\omega}\right).$
Equality holds if and only if one of the following holds:
  1. $(i)$
    If $\omega>2$, then $G$ is a regular $\omega$-partite graph.
  2. $(ii)$
    If $\omega=2$, then $G$ is a complete bipartite graph, and if $r$ is odd, then $G$ is regular.
When $r=1$, Theorem 1.18 implies Theorem 1.2 since $w_{1}(G)=n$. When $r=2$, Theorem 1.18 implies Theorem 1.3 since $w_{2}(G)=2m$. We obtain a localized strengthening of Nikiforov’s inequality.
Theorem 1.19.
Let $G$ be a graphFor any integer $r\geq 1$, we have
$\lambda_{1}^{r}(G)\leq\sum_{v\in V(G)}w_{r}(v)\sqrt{1-\frac{1}{c(v)}}\sqrt{1-\frac{1}{\omega(G)}}.$
The equality case is similar to that of Theorem 1.18.
Although Theorem 1.19 is localized, we believe that the following stronger localized version is true.
Conjecture 1.20.
For any integer $r\geq 1$, we have
$\lambda_{1}^{r}(G)\leq\sum_{v\in V(G)}w_{r}(v)\frac{c(v)-1}{c(v)}.$
1.4 Organization of the paper
This paper is organized as follows. In Section 2, we collect some known results about Motzkin-Straus type theorems, majorization, triangle-counting and random graphs, that will be useful later. Section 3 is devoted to Wilf-type results while in Section 4 we deal with generalizations of Bollobás-Nikiforov conjecture. Finally, in Section 5 we focus on inequalities involving walks.
2 Preliminaries
2.1 Weighted Motzkin-Straus type theorems
Consider a graph $G=(V(G),E(G))$ with $|V|$, the number of vertices, equals $n$. Let $S$ denote the standard simplex given by
$S=\left\{x\in\mathbb{R}^{n}:\sum_{v\in V(G)}x_{v}=1,x_{v}\geq 0,v\in V(G)\right\}.$
For a vector $x\in\mathbb{R}^{n}$, the support of $x$ is given by $\operatorname{supp}(x)=\{v\in V(G):x_{v}\neq 0\}$. Motzkin-Strauss [29] established the following remarkable result.
Theorem 2.1 ([29]).
Let $G$ be a graph with adjacency matrix $A(G)$ and clique number $\omega$. For any $x\in S$, we have
$x^{T}A(G)x\leq 1-\frac{1}{\omega}.$
Equality holds if and only if $\operatorname{supp}(x)$ induces a complete $\omega$-partite graph.
Instead of working with $A(G)$, one can also work with a weighted (non-negative weights) adjacency matrix $W(G)=[w_{uv}]\in\mathbb{R}^{n\times n}$ of the given graph $G$, and maximize the quantity $F_{G}(x):=x^{T}W(G)x$ on the simplex $S$. For the weight matrix $W(G)=[w_{uv}]$, where $w_{uv}=\frac{c(uv)}{c(uv)-1}$, it was shown in [7, 27]) that $F_{G}(x)\leq 1$ for all $x\in S$ and equality holds if and only if $\operatorname{supp}(x)$ induces a complete $\omega$-partite graph. Recently, Liu and Ning [26] established the following.
Lemma 2.2 ([26]).
Let $W(G)=[w_{uv}]$ be a weighted matrix where $w_{uv}=\frac{1}{2}\left(\frac{c(u)}{c(u)-1}+\frac{c(v)}{c(v)-1}\right)$. Then $F_{G}(x)\leq 1$ for all $x\in S$, and equality holds if and only if $\operatorname{supp}(x)$ induces a complete $\omega$-partite graph whose vertex classes $V_{1},\ldots,V_{\omega}$ satisfy $\sum_{v\in V_{i}}x_{v}=\frac{1}{\omega}$.
We observe the following corollary, which we require later.
Lemma 2.3.
Let $W(G)=[w_{uv}]$ be a weighted matrix where $w_{uv}=\sqrt{\frac{c(u)c(v)}{(c(u)-1)(c(v)-1)}}$. Then $F_{G}(x)\leq 1$ for all $x\in S$, and equality holds if and only if $\operatorname{supp}(x)$ induces a complete $\omega$-partite graph whose vertex classes $V_{1},\ldots,V_{\omega}$ satisfy $\sum_{v\in V_{i}}x_{v}=\frac{1}{\omega}$.
Proof.
Let $x\in S$. Using AM-GM inequality and Lemma 2.2, we get
$x^{T}W(G)x=\sum_{uv\in E(G)}2\sqrt{\frac{c(u)c(v)}{(c(u)-1)(c(v)-1)}}x_{u}x_{v}\leq\sum_{uv\in E(G)}\left(\frac{c(u)}{c(u)-1}+\frac{c(v)}{c(v)-1}\right)x_{u}x_{v}\leq 1.$
It is easily seen that the equality case is similar to that of Lemma $2.$ ∎
2.2 Majorization
For a vector $x\in\mathbb{R}^{n}$, let $x^{\downarrow}$ denote the vector obtained by rearranging the entries of $x$ in the non-increasing order. Given two vectors $x,y\in\mathbb{R}^{n}$, we say that $x$ is weakly majorized by $y$, if
$\sum_{j=1}^{k}{x_{j}}^{\downarrow}\leq\sum_{j=1}^{k}{y_{j}}^{\downarrow}$
for all $1\leq k\leq n$, and is denoted by $x\prec_{w}y$. Moreover, if $x\prec_{w}y\ \text{and}\ \sum_{i=1}^{n}x_{i}^{\downarrow}=\sum_{i=1}^{n}y_{i}^{\downarrow},$ then we say that $x$ is majorized by $y$, and is denoted by $x\prec y$. For $x=(x_{1},x_{2},\dots,x_{n})\in\mathbb{R}^{n}$ and $1\leq p<\infty$, define $\|x\|_{p}=\Big\{\sum_{i=1}^{n}|x_{i}|^{p}\Big\}^{\frac{1}{p}}.$
Theorem 2.4 ([24]).
Let $x=(x_{1},\dots,x_{n}),\ y=(y_{1},\dots,y_{n})\in\mathbb{R}^{n}_{\geq 0}$. If $y\prec_{w}x$, then
$\|y\|_{p}\leq\|x\|_{p}$
for any real number $p>1$, and equality holds if and only if $x=y$.
2.3 Triangle-counting
One way to count triangles in a graph is by counting the number of edges in vertex-neighbourhoods.
Lemma 2.5.
For every graph $G$, the number of triangles $t(G)$ in $G$ is given by
$t(G)=\frac{1}{3}\sum_{v\in V(G)}m(G[N(v)]),$
where $m(G[N(v)])$ denotes the number of edges in the subgraph induced by $N(v)$.
Let $F$ and $H$ be graphs. Let $\operatorname{ex}(n,H,F)$ denote the maximum number of copies of the graph $H$ in an $F$-free graph of order $n$. We mention the following general result concerning $F$-free graphs.
Theorem 2.6 ([16], Theorem 2).
Let $H$ and $F$ be graphs and assume $\chi(F)=k$. Then
$\operatorname{ex}(n,H,F)\leq\operatorname{ex}(n,H,K_{k})+o(n^{|V(H)|}).$
Let $\operatorname{mex}(m,H,F)$ denote the maximum number of copies of the graph $H$ in an $F$-free graph of size $m$. We state a similar counting result as above, but in terms of the size.
Theorem 2.7 ([33], Corollary 1.2).
Let $r\geq 3$ be an integer and $F$ be a graph. If $\operatorname{ex}(n,K_{r},F)=o(n^{s})$ for some $1<s\leq r$, then $\operatorname{mex}(m,K_{r},F)=o\left(m^{\frac{(r-1)s}{r+s-2}}\right)$.
We require the following corollary of the above theorems when $H$ is a triangle.
Corollary 2.8.
Let $F$ be a fixed graph with $\chi(F)=3$. Then
  1. $(i)$
    $\operatorname{ex}(n,K_{3},F)\leq o(n^{3})$
  2. $(ii)$
    $\operatorname{mex}(m,K_{3},F)\leq o(m^{3/2})$
Proof.
For $(i)$, taking $H=K_{3}$ in Theorem 2.6 and noting that $\operatorname{ex}(n,K_{3},K_{3})=0$, completes the proof. For $(ii)$, by the first assertion, we have $\operatorname{ex}(n,K_{3},F)\leq o(n^{3})$, implying that for some $\epsilon>0$, we have $\operatorname{ex}(n,K_{3},F)=o(n^{3-\epsilon})$. Now taking $r=3$ and $s=3-\epsilon$ in Theorem 2.7, gives $\operatorname{mex}(m,K_{3},F)=o\left(m^{\frac{6-2\epsilon}{4-\epsilon}}\right)\leq o(m^{3/2})$. ∎
We also require the following well-known result of Erdős and Gallai [12].
Theorem 2.9 (Erdős-Gallai [12]).
For a graph $G$ of order $n$ and size $m$ with no $P_{k}$ as a subgraph, we have
$m\leq\frac{n(k-2)}{2}.$
2.4 Random graphs
For $n\in\mathbb{N}$ and $0<p<1$, the Erdös-Renyi random graph $G(n,p)$ is a probability space over the set of graphs with vertex set $V(G)=\{v_{1},\ldots,v_{n}\}$ such that whenever $i<j$, the event that $v_{i}$ is adjacent to $v_{j}$ occurs with probability $p$, and these events are mutually independent.
Consider a symmetric matrix $A=[a_{ij}]$ of order $n$ such that $a_{ii}=0$ for all $i$ and $a_{ij}$’s are independent random variables whenever $i<j$ with probability mass function given by:
$\mathbb{P}(a_{ij}=1)=p\text{ and }\mathbb{P}(a_{ij}=0)=1-p,$
where $0<p<1$. Then $A$ is the adjacency matrix associated with the random graph $G(n,p)$. The eigenvalues of $G(n,p)$ are then defined to be the eigenvalues of $A$.
We say that a property $\Lambda$ holds for $G(n,p)$ aymptotically almost surely (a.a.s. for short) if
$\mathbb{P}(\Lambda\text{ holds for }G(n,p))\rightarrow 1\text{ as }n\rightarrow\infty.$
The following is known about the first and the second eigenvalue of random graphs.
Theorem 2.10.
For the random graph $G=G(n,p)$, the following hold $a.$s:
  1. $(i)$
    ([18]) $\lambda_{1}(G)=(p+o(1))n$.
  2. $(ii)$
    ([15]) $\lambda_{2}(G)\leq 2\left(\sqrt{p(1-p)}+o(1)\right)\sqrt{n}$.
The clique number of random graphs is given by the following well-known result.
Theorem 2.11 ([17, 5]).
For the random graph $G=G(n,p)$, a.a.s we have
$\omega(G)=\left(\frac{2}{\log(\frac{1}{p})}+o(1)\right)\log n.$
Using Wigner’s semicircle law [34], one can determine $s^{+}$ and $s^{-}$ of the random graph $G(n,1/2)$ as was done in [10].
Theorem 2.12 ([10]).
For the random graph $G=G(n,p)$, a.a.s we have
$s^{+}(G)=\frac{1}{2}\left(\frac{3}{4}+o(1)\right)n^{2}\text{ and }s^{-}(G)=\frac{1}{2}\left(\frac{1}{4}+o(1)\right)n^{2}.$
3 Wilf-type inequalities
3.1 Some general upper bounds
Firstly, we give an upper bound on $\sqrt{s^{+}}$ in terms of the number of triangles in the graph, thus proving Theorem 1.12.
Proof of Theorem 1.12.
It is well-known that
$\sum_{i=1}^{n}\lambda_{i}^{3}=\sum_{\lambda_{i}>0}\lambda_{i}^{3}-\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}=6t(G).$
We have
$s^{-}\geq\frac{\sum_{\lambda_{i}<0}|\lambda_{i}|^{3}}{|\lambda_{n}|}=\frac{\sum_{i=1}^{n^{+}}\lambda_{i}^{3}-6t(G)}{|\lambda_{n}|}\geq\frac{\sum_{i=1}^{n^{+}}\lambda_{i}^{3}-6t(G)}{\lambda_{1}}\geq\lambda_{1}^{2}-\frac{6t(G)}{\lambda_{1}},$
which gives
$\lambda_{1}\leq\sqrt{s^{-}+\frac{6t(G)}{\lambda_{1}}}.$ (1)
Thus,
$\sqrt{s^{+}}=\frac{\lambda_{1}}{\lambda_{1}}\sqrt{s^{+}}\leq\frac{1}{\lambda_{1}}\sqrt{s^{-}+\frac{6t(G)}{\lambda_{1}}}\sqrt{s^{+}}\leq\frac{1}{\lambda_{1}}\frac{2m+\frac{6t(G)}{\lambda_{1}}}{2}\leq\frac{n}{2}+\frac{3t(G)}{\lambda^{2}_{1}},$
proving the desired inequality.
It is clear from above that if equality holds, then $\lambda_{1}=|\lambda_{n}|$, implying $G$ is bipartite. Thus $s^{+}=s^{-}=m$. Now, $\sqrt{s^{+}}=\frac{n}{2}$ only if $G$ is complete bipartite. This completes the proof. ∎
We note here that inequality (1) in the above proof can be rearranged to obtain a lower bound on the number of triangles in a graph:
$t(G)\geq\frac{\lambda_{1}(G)(\lambda_{1}^{2}(G)-s^{-}(G))}{6}.$ (2)
One can compare (2) with the following well-known triangle counting result of Bollobás and Nikiforov [6]: for any graph $G$,
$t(G)\geq\frac{\lambda_{1}(G)(\lambda_{1}^{2}(G)-m)}{3}.$ (3)
Inequality (3) was also independently observed by Cioabă, Feng, Tait and Zhang [8], and the equality case was analyzed by Ning and Zhai [32]; also refer [23].
Owing to the difficulty of proving Conjecture 1.9 and 1.11, it is of interest to determine a non-trivial function $f:\mathbb{N}\rightarrow\mathbb{N}$ such that
$\sqrt{s^{+}(G)}\leq\sum_{v\in V(G)}\left(1-\frac{1}{f(c(v))}\right).$
We show that $f(\omega)=2\omega$ is sufficient.
Proposition 3.1.
For a graph $G$, we have
$\sqrt{s^{+}(G)}\leq\sum_{v\in V(G)}\left(1-\frac{1}{2c(v)}\right).$
Proof.
Let $y\in\mathbb{R}^{|V(G)|}$ be the vector such that $y_{v}=\sqrt{\frac{c(v)-1}{c(v)}}$ for all $v\in V(G)$. Now
$F_{G}(y)$ $=\sum_{uv\in E(G)}2\sqrt{\frac{c(u)c(v)}{(c(u)-1)(c(v)-1)}}y_{u}y_{v}=\sum_{uv\in E(G)}2=2m.$
Using Lemma 2.3, we have
$F_{G}(y)\leq\left(\sum_{v\in V(G)}y_{v}\right)^{2}=\left(\sum_{v\in V(G)}\sqrt{1-\frac{1}{c(v)}}\right)^{2}.$
Since $s^{+}+s^{-}=2m$, we have
$\sqrt{s^{+}}\leq\sqrt{2m}\leq\left(\sum_{v\in V(G)}\sqrt{1-\frac{1}{c(v)}}\ \right)\leq\sum_{v\in V(G)}\left(1-\frac{1}{2c(v)}\right).\qed$
For Conjecture 1.9, we can improve the function to $f(\omega)=2\omega-2$, thus proving Theorem 1.10.
Proof of Theorem 1.10.
For every vertex $v\in V(G)$, the subgraph induced by the open neighbourhood of $v$, i.e., $G[N(v)]$ is $K_{c(v)}$-free and has at most $\frac{\deg(v)^{2}}{2}\left(1-\frac{1}{c(v)-1}\right)$ edges by Turán’s Theorem. Using Lemma 2.5, we get
$t(G)\leq\sum_{v\in V}\frac{\deg(v)^{2}}{6}\left(1-\frac{1}{c(v)-1}\right).$ (4)
Since $\sum_{v\in V(G)}\deg(v)^{2}\leq n\lambda_{1}^{2}$ for any graph $G$, we get
$t(G)\leq\sum_{v\in V}\frac{\deg(v)^{2}}{6}\left(1-\frac{1}{\omega-1}\right)\leq\frac{n\lambda_{1}^{2}}{6}\left(1-\frac{1}{\omega-1}\right).$
Using Theorem 1.12, we see
$\sqrt{s^{+}}\leq\frac{n}{2}+\frac{3t(G)}{\lambda_{1}^{2}}\leq n\left(1-\frac{1}{2\omega-2}.\right)$
Next, we improve the above bound even further. Rearranging inequality 1 and noting that $s^{-}=2m-s^{+}$, we get
$s^{+}\leq 2m-\lambda_{1}^{2}+\frac{6t(G)}{\lambda_{1}}.$
Since $2m\leq n\lambda_{1}$ and substituting the bound for $t(G)$, we have
$s^{+}\leq n\lambda_{1}\left(2-\frac{1}{\omega-1}\right)-\lambda_{1}^{2}$
Wilf’s inequality gives us $\lambda_{1}\leq n\left(1-\frac{1}{\omega}\right)$, so maximing the function on the right gives
$s^{+}\leq n^{2}\left(1-\frac{1}{\omega}-\frac{1}{\omega^{2}}\right).$
This completes the proof. ∎
For regular graphs, the following is immediate from the proof of Theorem 1.10.
Proposition 3.2.
If $G$ is a regular graph, then
$\sqrt{s^{+}(G)}\leq\sum_{v\in V(G)}\left(1-\frac{1}{2c(v)-2}\right).$
3.2 Diamond-free graphs
Conjecture 1.11 is equivalent to Conjecture 1.9 for triangle-free graphs and hence true. We make further progress by proving Conjecture 1.11 for diamond-free graphs.
Theorem 3.3.
Let $G$ be a diamond-free graph of order $n\geq 42$. Then
$\sqrt{s^{+}(G)}\leq\sum_{v\in V(G)}\left(1-\frac{1}{c(v)}\right).$
Proof.
Observe that we can write the right-hand side of the inequality as follows:
$\sum_{v\in V(G)}\left(1-\frac{1}{c(v)}\right)=\frac{n}{2}+\sum_{v\in V(G)}\left(\frac{1}{2}-\frac{1}{c(v)}\right).$ (5)
Since $c(v)\geq 2$ for every vertex $v$, this gives that the right-hand side of the inequality in the assertion is always at least $n/2$.
If $G$ is triangle-free, then $t(G)=0$ and thus Theorem 1.12 implies $\sqrt{s^{+}}\leq\frac{n}{2}$. So, assume $t(G)\geq 1$. Since $G$ is diamond-free, it follows that $\omega(G)=3$ and $c(v)\leq 3$ for all vertices $v\in V$.
If $\lambda_{1}(G)\leq\frac{n}{4}$, then $\sqrt{s^{+}}\leq\sqrt{n\lambda_{1}}\leq\frac{n}{2}$, and we are done. So assume $\lambda_{1}\geq\frac{n}{4}$.
Since $G$ is diamond-free, for every vertex $v\in V(G)$, the induced subgraph $G[N(v)]$ is $P_{3}$-free and has at most $\frac{\deg(v)}{2}$ edges by Theorem 2.9. Using Lemma 2.5, we get
$t(G)\leq\sum_{v\in V(G)}\frac{\deg(v)}{2}=\frac{m}{3}.$
Applying Theorem 1.12 and using the fact that $2m\leq n\lambda_{1}$, we have
$\sqrt{s^{+}}\leq\frac{n}{2}+\frac{3t(G)}{\lambda^{2}_{1}}\leq\frac{n}{2}+\frac{m}{\lambda^{2}_{1}}\leq\frac{n}{2}+2.$ (6)
Define the set of triangular vertices in $G$ by
$TV(G)=\{v\in V(G):c(v)=3\},$
and let $tv(G)=|TV(G)|$. We consider the following cases:
Case 1: $tv(G)\geq 12$.
Then
$\sum_{v\in V(G)}\left(1-\frac{1}{c(v)}\right)\geq\frac{n}{2}+2,$
and the assertion holds by (6).
Case 2: $tv(G)\leq 11$.
Note that $3t(G)\leq{\frac{tv(G)(tv(G)-1)}{2}}\leq\frac{11\cdot 10}{2}$, implying $t(G)\leq\lfloor\frac{110}{6}\rfloor=18$. Using (6), we have
$\sqrt{s^{+}}\leq\frac{n}{2}+\frac{3t(G)}{\lambda^{2}_{1}}\leq\frac{n}{2}+\frac{3\cdot 18}{n^{2}/4^{2}}\leq\frac{n}{2}+\frac{1}{2},$
whenever $n\geq 42$. Since $t(G)\geq 1$, by (5) we have
$\sum_{v\in V(G)}\left(1-\frac{1}{c(v)}\right)\geq\frac{n}{2}+\frac{1}{2},$
completing the proof. ∎
3.3 $F$-free graphs
Theorem 3.4.
For any $F$-free graph $G$ with chromatic number $\chi(F)=3$, we have
$\sqrt{s^{+}(G)}\leq n\left(\frac{1}{2}+o(1)\right).$
Proof.
If $\lambda_{1}\leq\frac{n}{4}$, then $\sqrt{s^{+}}\leq\sqrt{n\lambda_{1}}\leq\frac{n}{2}$, and the assertion holds. So assume $\lambda_{1}\geq\frac{n}{4}$.
Using Corollary 2.8, we have $t(G)\leq o(n^{3})$. Using Theorem 1.12, we have
$\sqrt{s^{+}}\leq\frac{n}{2}+\frac{3t(G)}{\lambda^{2}_{1}}\leq\frac{n}{2}+\frac{o(n^{3})}{n^{2}}=n\left(\frac{1}{2}+o(1)\right).\qed$
3.4 Random graphs
Here, we prove Conjecture 1.11 for the random graph $G(n,\frac{1}{2})$.
Theorem 3.5.
For the random graph $G=G(n,\frac{1}{2})$, we have
$\sqrt{s^{+}(G)}\leq\sum_{v\in V(G)}\frac{c(v)-1}{c(v)}.$
Proof.
Note that for any vertex $v\in V(G)$, $\deg(v)=\left(\frac{1}{2}+o(1)\right)n$ $a.$s, and so
$c(v)=\omega(G[N(v)])=\omega(G(n/2,1/2))=\frac{2\log(n/2)}{\log(2)}$
$a.$s. Using Theorem 2.12, we see that $a.$s
$\sqrt{s^{+}}\leq n\sqrt{\frac{3}{8}+o(1)}\leq n\left(1-\frac{\log(2)}{2\log(n/2)}\right)=\sum_{v\in V(G)}\frac{c(v)-1}{c(v)}.\qed$
4 Generalizations of Bollobás-Nikiforov conjecture
4.1 Diamond-free graphs
In this section, we prove Conjecture 1.15 for diamond-free graphs, which is the next non-trivial case after triangle-free graphs. Kumar and Pragada [20] have shown that if diamond-free graphs with at least $34$ edges, then Conjecture 1.14 holds. We first show here that Conjecture 1.14 holds for all diamond-free graphs, as this is required later in the proof.
Theorem 4.1.
Let $G$ be a diamond-free graph. Then
$\lambda_{1}^{2}(G)+\lambda^{2}_{2}(G)\leq 2m\left(1-\frac{1}{\omega}\right).$
Proof.
We can assume that $n\geq 5$. Since $G$ is not a complete graph, $\lambda_{2}\geq 0$. Clearly, $\omega\leq 3$. The assertion is known to be true for triangle-free graphs, so we can assume that $\omega=3$. Suppose to the contrary that $\lambda_{1}^{2}+\lambda_{2}^{2}>\frac{4m}{3}$. Then,
$\sum_{\lambda_{i}<0}\lambda_{i}^{2}\leq 2m-(\lambda_{1}^{2}+\lambda_{2}^{2})<\frac{2m}{3}.$
Now, we count the number of triangles in $G$ in two different ways. As in the proof of Theorem 3.3, we have $t(G)\leq\frac{m}{3}$. Thus,
$2m\geq 6t(G)$ $=\sum_{i=1}^{n}\lambda_{i}^{3}$
$\geq\lambda_{1}^{3}+\lambda_{2}^{3}-\sum_{\lambda_{i}<0}|\lambda_{i}^{3}|$
$\geq\frac{(\lambda_{1}^{2}+\lambda_{2}^{2})^{3/2}}{\sqrt{2}}-\left(\sum_{\lambda_{i}<0}|\lambda_{i}^{2}|\right)^{3/2}$
$>\frac{1}{\sqrt{2}}\left(\frac{4m}{3}\right)^{3/2}-\left(\frac{2m}{3}\right)^{3/2}$
This is a contradiction whenever $m\geq 14$. If $m\leq 13$, then $n\leq 14$ since $G$ is connected, and the assertion is verified using a computer here. ∎
In [20], the following general bound was established for $\lambda_{1}^{2}+\lambda_{2}^{2}$ in terms of the number of triangles.
Theorem 4.2 ([20]).
Let $G$ be a graph of size $m$ and number of triangles $t(G)$. Then
$\lambda_{1}^{2}(G)+\lambda_{2}^{2}(G)<m+\left(3t(G)\right)^{2/3}.$
For diamond-free graphs, we improve the coefficient of $t(G)$ in the upper bound given in Theorem 4.2. Our proof relies on majorization, refer to Subsection 2.2.
Lemma 4.3.
Let $G$ be a diamond-free graph of size $m$ and number of triangles $t(G)$. Then
$\lambda_{1}^{2}(G)+\lambda_{2}^{2}(G)\leq m+\left(\frac{3}{\sqrt{2}}t(G)\right)^{2/3}.$
Proof.
We can assume that $n\geq 5$ and $\lambda_{2}\geq 0$. If $\lambda_{1}^{2}+\lambda_{2}^{2}\leq m$, then we are done. So suppose $\lambda_{1}^{2}+\lambda_{2}^{2}=m+\delta$, for some $\delta>0$. By Theorem 4.1, $\delta\leq m/3$, implying
$\lambda_{1}^{2}\geq\frac{\lambda_{1}^{2}+\lambda_{2}^{2}}{2}=\frac{m+\delta}{2}\geq 2\delta.$ (7)
Since $\sum_{i=1}^{n}\lambda_{i}^{2}=2m$, we have
$2(\lambda_{1}^{2}+\lambda_{2}^{2})=2m+2\delta=\sum_{i=1}^{n}\lambda_{i}^{2}+2\delta,$
which implies
$\lambda_{1}^{2}+\lambda_{2}^{2}=\sum_{i=3}^{n}\lambda_{i}^{2}+2\delta\geq\sum_{\lambda_{i}<0}^{n}\lambda_{i}^{2}+2\delta.$ (8)
Let $x=(\lambda_{1}^{2},\lambda_{2}^{2},0,\dots,0)^{T}$ and $y=(\lambda_{n}^{2},\lambda_{n-1}^{2},\dots,\lambda^{2}_{n-n^{-}+1},2\delta)^{T}$ in $\mathbb{R}^{n^{-}+1}$, where $n^{-}$ denotes the number of negative eigenvalues of $G$. Clearly, $\lambda_{1}^{2}\geq\max\{\lambda_{n}^{2},2\delta\}$ by (7). Then, using (8) we see that $y\prec_{w}x$. Applying Theorem 2.4 with $p=\frac{3}{2}$, we get $\|x\|^{3/2}_{3/2}\geq\|y\|^{3/2}_{3/2}$, that is,
$\lambda_{1}^{3}+\lambda_{2}^{3}\geq\sum_{\lambda_{i}<0}^{n}|\lambda_{i}|^{3}+(2\delta)^{3/2}.$
This implies that
$6t(G)$ $=\sum_{i=1}^{n}\lambda_{i}^{3}\geq\lambda_{1}^{3}+\lambda_{2}^{3}-\sum_{\lambda_{i}<0}^{n}|\lambda_{i}|^{3}$
$\geq(2\delta)^{3/2}=2\sqrt{2}\big(\lambda_{1}^{2}+\lambda_{2}^{2}-m\big)^{3/2}.$
Upon rearrangement, we get the desired inequality. ∎
Theorem 4.4.
Let $G$ be a diamond-free graph of size $m$ and number of triangles $t(G)$. If $t(G)\notin\{1,2,3,4\}$, then
$\lambda_{1}^{2}+\lambda_{2}^{2}\leq\sum_{e\in E(G)}2\left(1-\frac{1}{c(e)}\right).$
Proof.
If $t(G)=0$, then the assertion is known to be true. So assume $t(G)\geq 5$. Since $G$ is diamond-free, for any $e\in E(G)$, $c(e)\in\{2,3\}$. Moreover, if $c(e)=3$, then $e$ is contained in a unique triangle. Thus, we have
$\sum_{e\in E(G)}2\left(1-\frac{1}{c(e)}\right)=m+\sum_{e\in E(G)}\left(1-\frac{2}{c(e)}\right)=m+t(G).$ (9)
Since $t(G)\geq 5$ and using Lemma 4.3, we have
$\lambda_{1}^{2}+\lambda_{2}^{2}\leq m+\left(\frac{3}{\sqrt{2}}t(G)\right)^{2/3}\leq m+t(G).\qed$
When $t(G)\in\{1,2,3,4\}$, Lemma 4.3 gives an upper bound which is only slightly worse than $m+t(G)$.
4.2 $F$-free graphs
Theorem 4.5.
Let $F$ be a fixed graph with $\chi(F)=3$. Then every $F$-free graph $G$ of size $m$ satisfies
$\lambda_{1}^{2}(G)+\lambda_{2}^{2}(G)\leq m(1+o(1)).$
Proof.
Using Corollary 2.8, we have $t(G)\leq o(m^{3/2})$. Using Theorem 4.2, we get
$\lambda_{1}^{2}(G)+\lambda_{2}^{2}(G)\leq m+(3t(G))^{2/3}\leq m+o(m)=m(1+o(1)).\qed$
4.3 Random graphs
Here, we verify Conjecture 1.15 for the random graph $G(n,\frac{1}{2})$, thus showing that it is true for almost all graphs.
Theorem 4.6.
For the random graph $G=G(n,\frac{1}{2})$, we have
$\lambda_{1}^{2}(G)+\lambda_{2}^{2}(G)\leq\sum_{e\in E(G)}2\left(1-\frac{1}{c(e)}\right).$
Proof.
Observe that for any edge $uv\in E(G)$, we have
$c(uv)=\omega(G[N(u)\cap N(v)])=\omega(G(n/4,1/2))=\frac{2\log(n/4)}{\log 2}$
$a.$s. Using Theorem 2.12, we have $a.$s.
$s^{+}\leq n^{2}\left(\frac{3}{8}+o(1)\right)\leq\frac{n^{2}}{2}\left(1-\frac{\log 2}{2\log(n/4)}\right)=\sum_{e\in E(G)}2\left(1-\frac{1}{c(e)}\right).\qed$
5 Walk inequalities
In this section, we prove a localized version of Nikiforov’s Walk Inequality. First, we observe the following recursive inequality.
Lemma 5.1.
For any integer $r\geq 1$, we have
$w_{2r}(G)\leq\left(\sum_{v\in V(G)}w_{r}(v)\sqrt{\frac{c(v)-1}{c(v)}}\right)^{2}.$
Proof.
Let $y\in\mathbb{R}^{V(G)}$ be the vector such that $y_{v}=w_{r}(v)\sqrt{\frac{c(v)-1}{c(v)}}$ for all $v\in V(G)$. Now
$F_{G}(y)$ $=\sum_{uv\in E(G)}2\sqrt{\frac{c(u)c(v)}{(c(u)-1)(c(v)-1)}}y_{u}y_{v}=\sum_{uv\in E(G)}2{w_{r}(u)w_{r}(v)}=w_{2r}(G).$
Using Lemma 2.3, we have $F_{G}(y)\leq\left(\sum_{v\in V(G)}y_{v}\right)^{2}$ which proves the assertion. ∎
We are now ready to prove Theorem 1.19.
Proof of Theorem 1.19.
Using Nikiforov’s Walk Inequality (Theorem 1.18) and Lemma 5.1, we have
$\lambda_{1}^{2r}\leq\frac{\omega-1}{\omega}\,w_{2r}(G)\leq\frac{\omega-1}{\omega}\left(\sum_{v\in V(G)}w_{r}(v)\sqrt{\frac{c(v)-1}{c(v)}}\right)^{2}$
Taking square roots gives the desired result. Equality case is easily seen to be the same as in Theorem 1.18. ∎
Acknowledgement
M. Rajesh Kannan acknowledges financial support from the ANRF-CRG India and SRC, IIT Hyderabad. The authors thank Clive Elphick for several helpful comments. We also thank Rajat Adak, Yongtao Li and Xiao-Dong Zhang for suggesting some references.